# Sample complexity of optimal transport barycenters with discrete support

### v2

@@ 1 Introduction
Optimal transport barycenters provide a natural way for averaging a set of probability distributions while preserving their underlying geometric structure. They are formally defined as probability measures that minimize the following cost function
@@
where $\mathbb{D}$ denotes an optimal transport divergence, and $\left\{\mu^{\ell}\right\}_{\ell=1}^{L}$ is a collection of target probability measures in $\mathbb{R}^{d}$. This problem was first introduced and studied with the 2-Wasserstein metric squared $\mathbb{D}=W_{2}^{2}$ agueh2011barycenters, as a way to generalize McCann’s interpolation mccann1997convexity to more than two measures. Independently and in parallel, the authors of rabin2011wasserstein have proposed a method based on Wasserstein barycenters for texture synthesis and mixing. Then, following fast implementation of $W_{2}$-barycenters (e.g. cuturi2014fast), optimal transport barycenters became a central tool in machine learning and statistics. They have been used for various applications such as sample blending srivastava2015wasp, texture mixing rabin2011wasserstein, multisource domain adaptation montesuma2021wasserstein or music genre recognition montesuma2021wasserstein2. Although initially formulated with $\mathbb{D}=W_{2}^{2}$, other transport divergences have since been considered in (1.1), either to reduce their computational complexity, to induce desirable smoothing, or even to consider a more complex set of probability measures. This has led to the introduction of entropic chizat2023doubly, Bures (for centered Gaussian distributions) chewi2020gradient, sliced bonneel2015sliced and Gromov Wasserstein barycenters brogat2022learning.
Sample averaged approximation and barycenters
@@
There are many situations where the target probability measures in (1.1) may be intractable, either because they are modeled by continuous variables or because they are discretely supported on a very large number of point. To handle these cases from a computational point of view, one requires a discrete approximation. Typical schemes include histogram approximation (e.g. images), sample average approximation (SAA) montesuma2021wasserstein; claici2018stochastic or online stochastic approximation (SA) dvurechenskii2018decentralize; backhoff2024stochastic. This paper focuses on the sample average approximation of target measures, which primarily arises in two main use cases: the statistical setting where the target distributions are only accessible through empirical samples montesuma2021wasserstein and numerical monte-carlo approximation of known densities puccetti2020computation. In both cases, the targets are discrete empirical measures, and the resulting barycenter is therefore itself discrete. Consequently, most algorithms for computing optimal transport barycenters impose a constraint on the support size, similar to optimal quantization of measures pages2004optimal, which consists of approximating a probability measure with a small number of atoms. We include this support constraint in our treatement of the barycenter problem. In other words, we are interested in probability measures $\nu$ minimizing the function (1.1), subject to the constraint $|\operatorname{Supp}(\nu)|\leq N$ for some integer $N\geq 1$, where $\operatorname{Supp}(\nu)$ denotes the support of $\nu$. This problem is referred to as optimal transport barycenters with sparse support in the literature yang2024approximate; borgwardt2021computational. Many modern machine learning problems leverage this barycenter framework, including applications in LLM ai2025resmoe, flow cytometry datasets representation erell2024, graph clustering etienne2025pasco or cross-machine fault diagnosis lee2025barycenter. There exists few exceptions which do not include the support cardinality constraint; for example, staib2017parallel represents barycenters based on estimates of continuous functions by neural networks.
@@
In this article, we provide finite sample statistical guaranties for the sample averaged approximation of optimal transport barycenters under support cardinality constraints. The transport-based divergences $\mathbb{D}$ considered are the Wasserstein distances, the Sinkhorn divergences (entropic regularization), and the Sliced and max-Sliced Wasserstein distances. In other words, we consider a collection of target probability measures $\{\mu^{\ell}\}_{\ell=1}^{L}$ in $\mathbb{R}^{d}$, such that for each $\ell\in[\![1,L]\!]$, we have access to $n$ random variables $X_{1}^{\ell},\ldots,X_{n}^{\ell}\overset{\text{i.i.d}}{\sim}\mu^{\ell}$, and define their associated empirical measures by $\mu_{n}^{\ell}:=\frac{1}{n}\sum_{i=1}^{n}\delta_{X_{i}^{\ell}}$. The empirical sparse optimal transport barycenters are given by
@@
where $\mathbb{D}$ denotes any of the transport divergences stated above, $Y:=(y_{1},\ldots,y_{N})$ is a point cloud in $\mathbb{R}^{d}$ and $\Delta_{N}$ is the (N1)(N-1)-probability simplex.
Statistical results for optimal transport barycenters
In carlier2024quantitative authors study the statistical stability of empirical Wasserstein barycenter for general population measures (that is, measures supported on a set of probabilities), which includes the setting of (1.1). The rate of convergence they obtain suffers from the curse of dimensionality; this is inevitable in all generality as the barycenters inherit the sample complexity of optimal transport fournier2015rate. In heinemann2022randomized the authors consider the more specific framework of empirical barycenters of a collection of discrete target measures. Contrarily to carlier2024quantitative where the rate of convergence of empirical barycenter is $O(n^{-1/3d})$, the rate of convergence in this case is $O(C_{d}n^{-1/2})$, where the dimension of the target measures only appears in the constant $C_{d}$, and not exponentially in the sample size. As mentioned before, the fundamental difference in convergence rate behavior between the two frameworks lies in a similar phenomenon observed in optimal transport statistics, namely, that the statistical complexity of optimal transport adapts to the least complex measure hundrieser2024empirical. More precisely, in general, empirical optimal transport has a convergence rate of the form $O(n^{-1/d})$ fournier2015rate, while optimal transport between two probability measures, at least one of which is discrete, has a rate of convergence of $O(n^{-1/2})$ del2024central.
In the case of population barycenters, that is barycenters of a probability measure $\mathbb{Q}$ supported on the space of probability measures, numerous statistical results have been established bigot2020statistical; le2022fast. Although related, these results fall outside the scope of our study, as in their framework, the target measures $\mu^{\ell}$ are independently sampled from $\mathbb{Q}$. Let us note that in the context of entropy regularized Wasserstein barycenters, the sample average approximation was reported to be more efficient than the stochastic approximation approach in terms of computational complexity dvinskikh2020stochastic.
@@
We provide convergence bounds in expectation for the empirical sparse optimal transport barycenter functional (1.2) of the form $O(\sqrt{N/n})$, where $N$ denotes the number of support point of the barycenter, and $n$ is the number of samples per empirical target measure. These bounds are obtained for classical, entropy regularized, Sliced and max-Sliced Wasserstein barycenters, bridging the gap between various results in the literature. From this main result, we derive identical rates for the sample complexity of optimal transport barycenters with sparse support. Our proof techniques combine tools from empirical risk minimization theory and novel bounds on dual variables arising in the dual formulations of transport divergences.
@@
Throughout this article we will denote respectively $\mathcal{M}_{1}(\Omega)$ and $\mathcal{M}_{1}^{N}(\Omega)$, for some $\Omega\subset\mathbb{R}^{d}$, the set of probability measures on $\Omega$ and the set of probability measures supported on at most $N$ points in $\Omega$. We will denote $\mathbb{B}_{R}$ the closed ball of $\mathbb{R}^{d}$ centered in 0 and of radius $R>0$. Additionally, $\|\cdot\|$ is the Euclidean norm on $\mathbb{R}^{d}$, $\Delta_{N}$ the N1N-1 probability simplex and $\mathbb{S}^{d-1}$ the (d1)(d-1)-sphere in $\mathbb{R}^{d}$.
@@
We first introduce Wasserstein distances and divergences based on optimal transport, followed by optimal transport barycenters and their sparse approximation in Section 2. We then discuss their computation in Section 3. Our main results on sample complexity of barycenters are presented in Section 4, and the paper ends with a discussion in Section 5. Dual formulations of the divergences considered, together with technical details and proofs, are postponed to the appendices.
@@ 2.1 Optimal transport divergences
Optimal transport has become a powerful tool in machine learning due to its ability to compute distances between arbitrary probability measures (we refer the reader to peyre2019computational for a computational point of view, and to villani_opt_old_new for a more theoretical point of view). In mathematical terms, the optimal transport between two measures $\mu,\nu\in\mathcal{M}_{1}(\mathbb{R}^{d})$ consists in finding a coupling which minimizes the average costs of transporting $\mu$ onto $\nu$ :
@@
where $c:\mathbb{R}^{d}\times\mathbb{R}^{d}\to\mathbb{R}$ is a lower semi-continuous cost function and $\Pi(\mu,\nu)$ denotes the set of all probability couplings with marginals $\mu$ and $\nu$. When $c:=D^{p}$ is the $p$-exponent of a distance $D$, for $p\geq 1$, (2.1) defines the $p$-exponent of a distance $W_{p}$ in the space of probability measures, called the Wasserstein distance (villani_opt_old_new, Section 6), and is particularly suited as it metrizes the space of probability measures. In our case, we will consider the Euclidean distance on $\mathbb{R}^{d}$, and the Wasserstein distance is given by
@@
In the semi-discrete setting, where one measure is continuous and the other one is discrete, as considered in our sparse barycenter problem, this infinite dimensional optimization problem can be reformulated as a finite-dimensional maximization problem using the Kantorovich dual formulation of optimal transport (see Appendix A). Still, Wasserstein distances lack practicability as in the discrete setting for instance, solving eq. (2.2) involves addressing a potentially high-dimensional linear programming problem. This has motivated the introduction of new optimal transport-based divergences between probability measures.
The entropic regularization of Wasserstein distances, or Sinkhorn divergences cuturi2013sinkhorn; peyre2019computational, consists in adding an entropy regularization term to the optimal transport problem (2.1) and writes between two probability measures $\mu,\nu\in\mathcal{M}_{1}(\mathbb{R}^{d})$ as follows:
@@
where $\text{KL}(\gamma|\xi):=\int_{\mathbb{R}^{d^{2}}\times\mathbb{R}^{d^{2}}}(\log(\frac{d\gamma}{d\xi}(x,y))-1)d\gamma(x,y)$ denotes the Kullback-Leibler divergence and $\epsilon>0$ is a regularization parameter. The optimization problem in (2.3) can then be efficiently solved using Sinkhorn algorithm sinkhorn1967concerning; cuturi2013sinkhorn. The dual formulation of this entropy regularized optimal transport problem, studied in genevay2016stochastic with an emphasis on the discrete and semi-discrete settings, will be of particular use in our analysis and is presented in Appendix A. Note that it may be preferable to consider the debiased version of the Sinkhorn divergence genevay2018learning as entropy regularization can produce undesirable smoothing effect when computing sparse barycenters janati2020debiased. It is defined by
$\overline{W}_{\epsilon,p}^{p}(\mu,\nu)=W_{\epsilon,p}^{p}(\mu,\nu)-\frac{1}{2}(W_{\epsilon,p}^{p}(\mu,\mu)+W_{\epsilon,p}^{p}(\nu,\nu)).$ (2.4)
In any case, both classical and regularized optimal transport require solving an optimization problem as soon as the probability distributions lies in an Euclidean space of dimension greater than 1. For the univariate distributions case however, the Wasserstein distance (2.2) has an explicit expression which depends on the quantile function of the measures (ref_villani_topics, Chapter 2.2). By leveraging this analytic form through projections of the $d$-dimensional measures onto lines, rabin2011wasserstein have introduced sliced Wasserstein distances, which were then theoretically studied in kolouri2019generalized. More precisely, for any $\theta\in\mathbb{S}^{d-1}$, we set $P_{\theta}:x\in\mathbb{R}^{d}\mapsto\langle\theta,x\rangle\in\mathbb{R}$ the projection on the line directed by $\theta$. For $\mu,\nu$ probability measures in $\mathbb{R}^{d}$, one can then consider two collections of one-dimensional measures $(P_{\theta\sharp}\mu)_{\theta\in\mathbb{S}^{d-1}}$ and $(P_{\theta\sharp}\nu)_{\theta\in\mathbb{S}^{d-1}}$, where $T_{\sharp}\mu$ denotes the push forward (or image measure) of the measure $\mu$ by the measurable map $T$. For a fixed $\theta$, a sliced distance compares pairs $(P_{\theta\sharp}\mu,P_{\theta\sharp}\nu)$ of one-dimensional measures. If one averages these distances over all directions $\theta\in\mathbb{S}^{d-1}$, then the distance is called Sliced Wasserstein. If one chooses the direction that yields the larger distance between the projected measures, it is called max-Sliced Wasserstein. The Sliced and max-Sliced Wasserstein distance are therefore respectively defined as
@@
where $\sigma$ denotes the uniform probability measure over $\mathbb{S}^{d-1}$. Using the semi-dual formulation of $W_{p}$ one can also express these divergences in the semi-discrete case with a finite dimensional optimization program. These expressions are discussed in Appendix A.
@@ 2.2 Optimal transport barycenters with sparse support
An unconstrained barycenter of a collection of target probability measures $\{\mu^{\ell}\}_{\ell=1}^{L}$ with respect to an optimal transport divergence $\mathbb{D}$ is given as a minimizer of (1.1) over $\mathcal{M}_{1}(\mathbb{R}^{d})$. However, for computational reasons discussed in Section 3, this barycenter $\mu^{*}$ may be intractable in practice. For instance, for $\mathbb{D}=W_{p}^{p}$ with $p>1$, and if at least one of the target measures is absolutely continuous with respect to the Lebesgue measure, then the unique barycenter is also absolutely continuous agueh2011barycenters; brizzi2025p, and therefore inherently hard to recover. To address these situations, a common approach is to approximate $\mu^{*}$ by a discrete probability measure by restricting the optimization set in (1.1) to $\mathcal{M}_{1}^{N}(\mathbb{R}^{d})$. This approach is especially relevant when the target measures are available in discretized form (e.g, as empirical measures in an SAA framework), since the resulting barycenter is naturally discrete as well, see Section 5.2. In the literature, this is referred to as optimal transport barycenter problem with sparse support, or equivalently, discretely supported barycenters with a cardinality constraint. Formally, the aim is to find a probability measure $\nu^{*}:=\sum_{i=1}^{N}\pi_{i}\delta_{y_{i}}$ where $Y:=(y_{1},\ldots,y_{N})\in(\mathbb{R}^{d})^{N}$ and $\pi\in\Delta_{N}$ , that solves the following optimization problem:
@@
where A is a closed nonempty subset of $(\mathbb{R}^{d})^{N}\times\Delta_{N}$. When specifying the set of optimization constraint $A$, and the number of measures $L$, one recovers various existing cases in the literature cuturi2014fast; borgwardt2021computational that we summarize in the following Table 1.
@@
Table 1: Classification of problem (2.6) depending on the constraint set AA and the number of measures $L$.
In this article, we provide statistical complexity bounds for (2.6) with $\mathbb{D}=W_{p}^{p},\>W_{\epsilon,p}^{p},\>\overline{W}_{\epsilon,p}^{p},\>SW_{p}^{p}$ and $\text{max-}SW_{p}^{p}$.
3 Algorithmic aspects of optimal transport barycenters and their sparse approximation
Computing optimal transport barycenters presents two main challenges: the need to discretize the target measures and the inherent computational complexity of optimal transport.
Discretization of target measures
As mentioned before, although the target probability distributions may be continuous in the optimization problem (1.1), an algorithm to compute the barycenter operates on discrete data, as it can only handle finite objects. Generally, continuous target probability measures are therefore approximated by empirical measures, and the barycenter (1.1) is computed either by (deterministic) gradient or fixed-point methods cuturi2014fast, or by stochastic gradient methods li2020continuous. Similarly, for discrete target distributions supported on a large number of atoms, one typically resorts to empirical approximations of the whole distributions and stochastic gradient methods heinemann2022randomized.
Additionally, most practical algorithms implicitly constrain the cardinality of the barycenter, whose number of atoms must be specified as input. Therefore, these algorithms actually solve the constrained problem (2.6) instead of (1.1). This has led to the notion of optimal transport barycenters with sparse support yang2024approximate; borgwardt2021computational. This framework was recently considered in mimouni2025computation, in which the authors propose an algorithm to solve (2.6) for empirical measures. Note that these barycenters are strongly related to optimal quantization and K-means problems (consider $L=1$ in (2.6)), as discussed in Section (5.4).
For (centered) Gaussian target distributions however, a fixed-point algorithm on covariance matrices alvarez2016fixed allows to compute the Gaussian barycenter agueh2011barycenters without discretization. Finally, an alternative to directly handling continuous measures could be to triangulate the supports of the target distributions, as is done in the optimal quantization problem and semi-discrete optimal transport levy2015numerical. However, this approach seems difficult to extend to the barycenter problem since target measures may be distributed over very different supports.
Computational complexity of optimal transport.
Even for discrete probability distributions supported on $N$ points in $\mathbb{R}^{d}$, the optimal transport problem has a prohibitive $O(N^{3}\log(N))$ computational cost peyre2019computational. The Wasserstein barycenter problem naturally inherits this complexity, for a set of $L$ measures altschuler2022wasserstein; borgwardt2021computational. Consequently, the barycenter problem is only tractable for measures supported on a reasonable number of atoms altschuler2021wasserstein.
In practice, Sinkhorn divergences $W_{\epsilon,p}^{p}$ (2.3) are often favored for computing Wasserstein barycenters (1.2), resulting in Sinkhorn barycenters. These divergences offer a significant computational advantage, of order $O(N^{2}/\epsilon)$dvurechensky2018computational, which allows a trade-off between optimal transport accuracy and algorithmic efficiency. Regularization is crucial and can be introduced in two ways : either by using Sinkhorn divergence $\mathbb{D}=W_{\epsilon,p}^{p}$ in the barycenter problem, or by adding an entropic regularization term on the barycenter itself in the optimization problem (1.2). In chizat2023doubly, the author studies both regularization approaches, as well as a combination of the two.
Sliced-Wasserstein distances are also relevant to compute barycenters in the space of probability measures, as introduced in rabin2011wasserstein for an application in texture mixing. The Sliced-Wasserstein barycenters were then studied in bonneel2015sliced. As mentioned in Section 2, Sliced distances rely on the closed-form solution of the one-dimensional Wasserstein distance, and Sliced barycenters consequently inherit this computational efficiency.

### v3

@@ 1 Introduction
Optimal transport barycenters provide a geometrically structured notion of mean for a set of probability measures. They are minimizers of functionals of the form
@@
where $\nu$ is a probability measure, $\mathbb{D}$ is an optimal transport divergence, and $\left\{\mu^{\ell}\right\}_{\ell=1}^{L}$ is a collection of target probability measures in $\mathbb{R}^{d}$. This problem was first introduced for $\mathbb{D}=W_{2}^{2}$ agueh2011barycenters, as a generalization of McCann’s interpolation mccann1997convexity. This was also proposed independently for texture synthesis and mixing rabin2011wasserstein. Following fast implementation of $W_{2}$-barycenters (e.g. cuturi2014fast), they became a central tool in machine learning and statistics with various applications such as sample blending srivastava2015wasp, texture mixing rabin2011wasserstein, multisource domain adaptation montesuma2021wasserstein or music genre recognition montesuma2021wasserstein2. Various extensions based on different transport divergences $\mathbb{D}$ in (1.1) exist: entropic chizat2023doubly, Bures chewi2020gradient, sliced bonneel2015sliced and Gromov Wasserstein barycenters brogat2022learning.
Sample averaged approximation and sparse barycenters
@@
In many applications, the target probability measures in (1.1) are computationally intractable, and one needs to resort to discrete approximations. Typical schemes include histogram approximation (e.g. images), sample average approximation (SAA) montesuma2021wasserstein; claici2018stochastic or online stochastic approximation (SA) dvurechenskii2018decentralize; dvinskikh2020stochastic; backhoff2024stochastic. We focus on sample average approximation arising in two main use cases: the statistical setting where distributions are only accessible through empirical samples montesuma2021wasserstein and numerical Monte-Carlo approximation of known densities puccetti2020computation. In both cases, all measures, including the resulting barycenter, are discrete, possibly supported on a very large number of atoms. Consequently, most optimal transport barycenter algorithms restrict the number of atoms of the barycenter, similarly to optimal quantization of measures pages2004optimal. Our analysis includes this support constraint. In other words, the probability measures $\nu$ in (1.1), is subject to the constraint $|\operatorname{Supp}(\nu)|\leq N$ for some integer $N\geq 1$, where $\operatorname{Supp}(\nu)$ denotes the support of $\nu$. This problem is known as optimal transport barycenters with sparse support yang2024approximate; borgwardt2021computational. Many modern machine learning problems leverage this sparse barycenter framework, including applications in LLM ai2025resmoe, flow cytometry erell2024, graph clustering etienne2025pasco or cross-machine fault diagnosis lee2025barycenter.
@@
Our main problem of interest is that of statistical guaranties for the optimal transport barycenters sample averaged approximation under support cardinality constraints, with various divergences $\mathbb{D}$: (sliced) Wasserstein distances, (debiased) Sinkhorn divergences. The collection of target probability measures $\{\mu^{\ell}\}_{\ell=1}^{L}$ in $\mathbb{R}^{d}$ each relate to $n$ i.i.d random variables $X_{1}^{\ell},\ldots,X_{n}^{\ell}\overset{\text{i.i.d}}{\sim}\mu^{\ell}$ with empirical measures by $\mu_{n}^{\ell}:=\frac{1}{n}\sum_{i=1}^{n}\delta_{X_{i}^{\ell}}$, and the empirical sparse barycenters are given by
@@
where $Y:=(y_{1},\ldots,y_{N})$ is a point cloud in $\mathbb{R}^{d}$ and $\Delta_{N}$ is the (N1)(N-1)-probability simplex.
Statistical guaranties for optimal transport barycenters
For general population measures, statistical rates of the form $O(n^{-1/3d})$ for empirical barycenters were described in carlier2024quantitative, suffering from the curse of dimensionality. On the other hand, for discrete target measures, an $O(n^{-1/2})$ rate is known heinemann2022randomized. This is reminiscent of the statistical complexity of optimal transport adapting to the least complex measure hundrieser2024empirical. Indeed, empirical optimal transport converges as $O(n^{-1/d})$ fournier2015rate, while if at least one of the two measures is discrete, the rate is $O(n^{-1/2})$ del2024central.
@@
We provide uniform rates in expectation for the empirical sparse optimal transport barycenter functional (1.2). They are of the form $O(\sqrt{N/n})$, where the barycenter has $N$ support points, and $n$ is the number of samples per empirical target measure. We consider several divergences bridging the gap between various results in the literature. We deduce sample complexity estimates for the optimal transport barycenters with sparse support. Our proof techniques combine tools from empirical risk minimization theory and estimations on dual variables arising in the dual formulations of transport divergences.
@@
Throughout this article we will denote respectively $\mathcal{M}_{1}(\Omega)$ and $\mathcal{M}_{1}^{N}(\Omega)$, for some $\Omega\subset\mathbb{R}^{d}$, the set of probability measures on $\Omega$ and the set of probability measures supported on at most $N$ points in $\Omega$. We will denote $\mathbb{B}_{R}$ the closed ball of $\mathbb{R}^{d}$ centered in 0 and of radius $R>0$. Additionally, $\|\cdot\|$ is the Euclidean norm on $\mathbb{R}^{d}$, $\Delta_{N}$ the N1N-1 probability simplex and $\mathbb{S}^{d-1}$ the (d1)(d-1)-sphere in $\mathbb{R}^{d}$. For a measure $\mu$ and a measurable map $T$ we denote by $T_{\sharp}\mu$ the pushforward (or image measure) of $\mu$ by $T$.
@@
We introduce technical material on optimal transport divergences in Section 2. Computational aspects are discussed in Section 2.3. Our main results are presented in Section 3, with a discussion in Section 4. Dual formulations of the divergences considered, together with technical details and proofs, are postponed to the appendices.
@@ 2.1 Optimal transport divergences
We refer to peyre2019computational for a computational point of view, and to villani_opt_old_new for a more theoretical presentation. The optimal transport between two measures $\mu,\nu\in\mathcal{M}_{1}(\mathbb{R}^{d})$ is a solution to the optimization problem over measures:
@@
where $c:\mathbb{R}^{d}\times\mathbb{R}^{d}\to\mathbb{R}$ is a lower semi-continuous cost function and $\Pi(\mu,\nu)$ is the set of couplings: probability measures with marginals $\mu$ and $\nu$. In the Euclidean case, if $c\colon(x,y)\mapsto\|x-y\|^{p}$ for an integer $p\geq 1$, this is the Wasserstein distance $W_{p}$ on $\mathbb{R}^{d}$:
@@
The semi-discrete setting corresponds to one measure being finitely supported. We focus on this setting which encompasses the sparse barycenter problem. The infinite dimensional problem over measures can be reformulated as a finite-dimensional maximization problem based on optimal transport duality (see Appendix A). Various optimal transport-based divergences were introduced based on variations of the Wasserstein distance.
The Sinkhorn divergences cuturi2013sinkhorn; peyre2019computational add an entropic regularization term, for $\mu,\nu\in\mathcal{M}_{1}(\mathbb{R}^{d})$:
@@
where $\text{KL}(\gamma|\xi):=\int_{\mathbb{R}^{2d}\times\mathbb{R}^{2d}}(\log(\frac{d\gamma}{d\xi}(x,y))-1)d\gamma(x,y)$ denotes the Kullback-Leibler divergence and $\epsilon>0$ is a regularization parameter. For discrete measures, the optimization problem in (2.3) can then be efficiently solved using the Sinkhorn algorithm sinkhorn1967concerning; cuturi2013sinkhorn. Similarly as Wasserstein distances, in the semi-discrete setting, there is a finite dimensional dual problem genevay2016stochastic of particular use in our analysis, it is presented in Appendix A.
Considering projections onto lines leads to sliced Wasserstein distances rabin2011wasserstein; kolouri2019generalized. For any $\theta\in\mathbb{S}^{d-1}$, set $P_{\theta}:x\in\mathbb{R}^{d}\mapsto\langle\theta,x\rangle\in\mathbb{R}$, understood as the projection on the line directed by $\theta$. For $\mu,\nu\in\mathcal{M}_{1}(\mathbb{R}^{d})$, one obtains two collections of pushforward univariate measures $(P_{\theta\sharp}\mu)_{\theta\in\mathbb{S}^{d-1}}$ and $(P_{\theta\sharp}\nu)_{\theta\in\mathbb{S}^{d-1}}$. Sliced distances combine divergences between these univariate measures for all directions $\theta\in\mathbb{S}^{d-1}$. We obtain the Sliced and max-Sliced Wasserstein distances respectively defined as
@@
where $\sigma$ denotes the uniform probability measure over $\mathbb{S}^{d-1}$. In the semi-discrete setting, these also admits finite dimensional dual problems as discussed in Appendix A.
@@ 2.2 Optimal transport barycenters with sparse support
Given a collection of target probability measures $\{\mu^{\ell}\}_{\ell=1}^{L}$, an unconstrained barycenter with respect to divergence $\mathbb{D}$ is a minimizer of (1.1), it is typically absolutely continuous agueh2011barycenters; brizzi2025p if at least one of the target measures is absolutely continuous. A common strategy is approximation by discrete probability measures, restricting the optimization set in (1.1) to $\mathcal{M}_{1}^{N}(\mathbb{R}^{d})$. This is referred to as optimal transport barycenter problem with sparse support, with underlying measures of the form $\nu^{*}:=\sum_{i=1}^{N}\pi_{i}\delta_{y_{i}}$ where $Y:=(y_{1},\ldots,y_{N})\in(\mathbb{R}^{d})^{N}$ and $\pi\in\Delta_{N}$. The problem is reformulated as follows:
@@
where A is a closed nonempty subset of $(\mathbb{R}^{d})^{N}\times\Delta_{N}$ which integrates possibly additional constraints. Formulation (2.5) covers various settings described in the literature cuturi2014fast; borgwardt2021computational:
@@
2.3 Discretization and computational complexity
Even for continuous target measures, barycenter algorithms operate on discrete data. A common approach is to approximate the target measures by empirical averages, and to compute their barycenter (1.1) with dedicated methods cuturi2014fast; li2020continuous; heinemann2022randomized. Most algorithms actually add support constraints as in (2.5), effectively computing a sparse barycenters yang2024approximate; borgwardt2021computational; mimouni2025computation. For discrete measures supported on $N$, the optimal transport problem has an $O(N^{3}\log(N))$ computational cost peyre2019computational. The discrete Wasserstein barycenter problem inherits this complexity altschuler2022wasserstein; borgwardt2021computational, making it tractable only for measures supported on a small number of atoms altschuler2021wasserstein. The complexity of Sinkhorn divergences $W_{\epsilon,p}^{p}$ (2.3) is $O(N^{2}/\epsilon)$dvurechensky2018computational, which allows for a trade-off between optimal transport accuracy and algorithmic efficiency chizat2023doubly. Sliced-Wasserstein distances are based on univariate closed form quantile expressions (ref_villani_topics, Chapter 2.2), making them also computationally relevant bonneel2015sliced.